# Dynamical mean field approach to associative memory model with non-monotonic transfer functions

### v1

@@ IV.1 Comparison between DMFT and direct Simulations
We compared the results obtained from the DMFT effective dynamics Eq. (32) with those from direct simulations of Eq. (11). For the neuronal transfer function, we used Eq. (8) with the parameter set $(c,c^{\prime},h,k)=(50,15,0.5,-0.5)$, as specified in Morita’s original work [7]. The system sizes were set to $N_{\rm DMFT}=10^{6}$ and $N_{\rm direct}=4096$ for the DMFT and direct simulations, respectively, and the time step was fixed at $\gamma=0.1$.
@@
is shown in Fig. 3. Except for the critical cases, excellent agreement is observed in both figures, confirming that the DMFT approach reliably captures the macroscopic dynamical behavior of large systems.
@@ IV.2 Assessment of self-feedback effects
Beyond enabling access to the macroscopic properties of large systems, another advantage of the DMFT framework is its ability to evaluate self-feedback effects explicitly. Figure 5 plots the feedback coefficients $\Lambda(t,s)$ for both successful [panels (a), (b)] and failed [panels (c), (d)] retrieval cases. When retrieval succeeds, $\Lambda(t,s)$ takes non-negligible values once the system has macroscopically converged to the retrieval state (around $t\gtrsim 20$, see Figs. 3 and 4) and continues to exhibit irregular oscillations thereafter. Although the retrieval state appears macroscopically stationary and seems to satisfy time-translation invariance, $\Lambda(t,s)$ itself is not stationary.
@@
This apparent irregularity is, however, not intrinsic to the dynamics itself but can be attributed to Monte Carlo errors arising from the stochastic nature of the Langevin updates. When the system converges to a retrieval state, the random variable $\phi_{i}^{t}$—originating from the cross-talk noise $N^{-1/2}\sum_{\mu=2}^{p}\xi_{i}^{\mu}\sum_{j=1}^{N}\xi_{j}^{\mu}f(a_{j}^{t})$—also converges to a constant. Consequently, the noise correlation approaches a finite value in the long-time limit, $C(t,s)\to\sigma_{\phi}^{2}$ as $t,s\to\infty$. Similarly, $K(t,s)=N^{-1}\sum_{i=1}^{N}f(a_{i}^{t})\phi_{i}^{s}$ tends to a constant $K^{*}$ in the same limit. This implies that, for large $t$, the matrix $C_{t-1}$ possesses very small eigenvalues in directions orthogonal to $(1,\ldots,1)^{\top}$. As a result, when $C_{t-1}^{-1}$ acts in Eq. (19), Monte Carlo fluctuations along those orthogonal directions in $G_{t}(t,s)$ are strongly amplified, which in turn leads to large fluctuations in $\Lambda_{t}(t,s)$ through Eq. (20).
For the asymptotic regime where $t,s\gg 1$, the component of $\Lambda(t,s)$ (which is equal to $\Lambda_{t}(t,s)$) with the fluctuation term removed is evaluated as follows. In this context, it is reasonable to assume $C(t,s)=\sigma_{\phi}^{2}$ and $K(t,s)=K^{*}$ for all pairs of $t$ and $s$. The component of $G(t,s)$ in the direction of $(1,\ldots,1)^{\top}$ is then derived from equation (19) as:
@@
The behavior of Eq. (35) depends on the sign of $K^{*}$. When $K^{*}>0$ the feedback coeffcient $\Lambda(t,s)$ of past states at time $s<t$ on a given time $t$ becomes stronger as one goes further back in time. In contrast, when $K^{*}<0$, the most recent past exerts the strongest effect, and the influence gradually decays as one traces further back in time. However, in either case, individual feedback components from past times $s$ vanish algebraically as $t$ increases [Fig. 7 (a,b)].
@@
for $t\gg 1$ when $\sigma_{\phi}^{-2}K^{*}<1$, which we assume here. Furthermore, since the fluctuations are orthogonal to $(1,\ldots,1)^{\top}$, they are expected to cancel out in Eq. (38). Indeed, the behavior of $\Lambda(t)$ observed in successful retrieval cases supports this scenario [Fig. 8(a)]. Conversely, when retrieval fails, $C(t,s)$ decays as |ts||t-s| increases [Fig. 6(c,d)], and the feedback coefficients $\Lambda(t,s)$ remain small regardless of $t$ and $s$, while its itegrated form continues to move slowly [Fig. 8(b)]. This behavior suggests that the system remains in a chaotic state.
@@ IV.3 Connection to earlier studies
for each $x$ (see Appendix C for details). The convergent value of the readout overlap is assessed as $M=N_{\rm DMFT}^{-1}\sum_{i=1}^{N_{\rm DMFT}}{\rm sign}(m+\phi_{i})$. Comparison of Eqs. (41) and (45) indicates that $U=\sigma_{\phi}^{-2}K^{*}$ holds. Eqs. (42)–(46) are equivalent to the self-consistent signal-to-noise analysis (SCSNA) developed for analog neural networks in Ref. [18], provided that the sample averages $N_{\rm DMFT}^{-1}\sum_{i=1}^{N_{\rm DMFT}}(\cdots)$ are replaced by Gaussian averages $\int d\phi\,\mathcal{N}(\phi|0,\sigma_{\phi}^{2})(\cdots)$, which is justified by the law of large numbers in the limit of $N_{\rm DMFT}\to\infty$. Similar results can also be found in [19].
Ref. [18] shows that SCSNA can accurately describe macroscopic quantities such as storage capacity and the local-field distribution as long as the function $g(x)$ is uniquely determined by Eq. (46). However, it also reports that when multiple solutions of Eq. (46) emerge as the profile of $f(a)$ is gradually modified, the SCSNA predictions become unreliable. Indeed, in the present model, the convergent DMFT solution reveals that Eq. (46) admits multiple branches [Fig. 9]. For each input $m+\phi_{i}$ in Eqs. (42)–(44), the appropriate branch of $g(m+\phi_{i})$ is determined by the entire history of the DMFT dynamics. This implies that DMFT is indispensable for examining the fixed-point structure of the retrieval dynamics.
Figure 9: Multiple branches of $g(x)$. Markers represent 1000 randomly selected elements at the 1000th iteration of Eq. (32), at which DMFT is considered to have practically converged.
@@ Acknowledgements.
Let us denote $(T+1)\times(T+1)$ matrices composed of $Q(t,s)$, $R(t,s)$, and $S(t,s)$ in Eqs. (21)-(23) by $Q$, $R$, and $S$, respectively. Using these quantities, we define a $2(T+1)\times 2(T+1)$ matrix as
@@
and similarly for $R(t,s)$ and $S(t,s)$ for $\forall{t},\forall{s}\in\{0,\ldots,T\}$, into the right-hand side of Eq. (18). Evaluating integrals over $Q$, $R$, and $S$ using the saddle point method yields
@@
$\times\!\exp\!\left(-\!\sum_{i=1}^{N}\!\sum_{t,s}\!\left[\tfrac{1}{2}\hat{R}(t,s)(ih_{i}^{t})(ih_{i}^{s})+\hat{S}(t,s)f(a_{i}^{t})(ih_{i}^{s})\right]\!\right)\!\left.\times\exp\!\left(\sum_{t=0}^{T+1}\bm{\theta}^{t}\!\cdot\!\bm{a}^{t}\right)\right\}$ (58)
$=$ $\,\mathop{\rm tr}_{\bm{\sigma}^{0},\tilde{\bm{a}},\tilde{\bm{h}},\tilde{\bm{\phi}}}\!\left\{P(\bm{\sigma}^{0})\exp\!\left[-\!\sum_{t=0}^{T}\sum_{i=1}^{N}ih_{i}^{t}{\mathcal{L}}_{i}^{t}\right]\times\!\exp\!\left(\bm{\theta}^{t}\!\cdot\!\bm{a}^{t}\right)\right$. (59)
@@
$\mathbb{E}_{\tilde{\bm{\phi}}}\left[\frac{\partial f(a_{i}^{t})}{\partial\phi_{i}^{s}}\right]=\mathbb{E}_{\tilde{\bm{\phi}}}\left[f(a_{i}^{t})\left(\sum_{w}\phi_{i}^{w}\gamma^{2}\hat{R}^{-1}(w,s)\right)\right],$ (67)

### v2

@@ IV.1 Comparison between DMFT and direct Simulations
We compared the results obtained from the DMFT effective dynamics Eq. (29) with those from direct simulations of Eq. (11). For the neuronal transfer function, we used Eq. (8) with the parameter set $(c,c^{\prime},h,k)=(50,15,0.5,-0.5)$, as specified in Morita’s original work [7]. The system sizes were set to $N_{\rm DMFT}=10^{6}$ and $N_{\rm direct}=4096$ for the DMFT and direct simulations, respectively, and the time step size was fixed at $\gamma=0.1$.
@@
is shown in figure 3. Except for the critical cases, excellent agreement is observed in both figures, confirming that the DMFT approach reliably captures the macroscopic dynamical behavior of large systems.
@@
The expansion of the basin of attraction induced by non-monotonic transfer functions can be understood as follows [8]. Increasing the number of stored patterns enlarges the variance of the cross-talk noise. As a result, an increasing number of neurons receive large inputs with the opposite sign from that of the target pattern to be retrieved. When a monotonic transfer function is used, the presence of many such neurons that are “overconfident” yet incorrect undermines the stability of the retrieval state. In contrast, with a non-monotonic transfer function, neurons receiving excessively large inputs flip their output. Consequently, such overconfident neurons cannot persist, which makes the variance of the cross-talk noise remain relatively small compared with the signal components that convey information about the target pattern. This allows the retrieval state to remain stable up to larger pattern ratios $\alpha$ than in the case with a monotonic transfer function.
@@ IV.2 Assessment of self-feedback effects
Beyond enabling access to the macroscopic properties of large systems, another advantage of the DMFT framework is its ability to evaluate self-feedback effects explicitly. Figure 5 plots the feedback coefficients $\Lambda(t,s)$ for both successful (panels (a), (b)) and failed (panels (c), (d)) retrieval cases. When retrieval succeeds, $\Lambda(t,s)$ takes non-negligible values once the system has macroscopically converged to the retrieval state (around $t\gtrsim 20$, see figures 3 and 4) and continues to exhibit irregular oscillations thereafter. Although the retrieval state appears macroscopically stationary and seems to satisfy time-translation invariance, $\Lambda(t,s)$ itself is not stationary.
@@
This apparent irregularity is, however, not intrinsic to the dynamics itself but can be attributed to Monte Carlo errors arising from the stochastic nature of the Langevin updates. When the system converges to a retrieval state, the random variable $\phi_{i}^{t}$—originating from the cross-talk noise $N^{-1/2}\sum_{\mu=2}^{p}\xi_{i}^{\mu}\sum_{j=1}^{N}\xi_{j}^{\mu}f(a_{j}^{t})$—also converges to a constant. Consequently, the noise correlation approaches a finite value in the long-time limit (figures 6 (a,b)), $C(t,s)\to\sigma_{\phi}^{2}$ as $t,s\to\infty$, and the correlation between $f(a_{i}^{t})$ and $\phi_{i}^{t}$, $K(t,s)=N^{-1}\sum_{i=1}^{N}f(a_{i}^{t})\phi_{i}^{s}$, which appears in the right-hand side in Eq. (32), also tends to a constant $K^{*}$ in the same limit. This implies that, for large $t$, the matrix $C_{t-1}$ possesses very small eigenvalues in directions orthogonal to $(1,\ldots,1)^{\top}$. As a result, when $C_{t-1}^{-1}$ acts in Eq. (32), Monte Carlo fluctuations along those orthogonal directions in $G_{t}(t,s)$ are strongly amplified, which in turn leads to large fluctuations in $\Lambda_{t}(t,s)$ through Eq. (33).
For the asymptotic regime where $t,s\gg 1$, the component of $\Lambda(t,s)$ (which is equal to $\Lambda_{t}(t,s)$) with the fluctuation term removed is evaluated as follows. In this context, it is reasonable to assume $C(t,s)=\sigma_{\phi}^{2}$ and $K(t,s)=K^{*}$ for all pairs of $t$ and $s$. The component of $G(t,s)$ in the direction of $(1,\ldots,1)^{\top}$ is then derived from Eq. (32) as:
@@
The behavior of Eq. (35) depends on the sign of $K^{*}$. When $K^{*}>0$ the feedback coeffcient $\Lambda(t,s)$ of past states at time $s<t$ on a given time $t$ becomes stronger as one goes further back in time. In contrast, when $K^{*}<0$, the most recent past exerts the strongest effect, and the influence gradually decays as one traces further back in time. However, in either case, individual feedback components from past times $s$ vanish algebraically as $t$ increases (figures 7 (a,b)).
@@
for $t\gg 1$ when $\sigma_{\phi}^{-2}K^{*}<1$, which we assume here. Furthermore, since the fluctuations are orthogonal to $(1,\ldots,1)^{\top}$, they are expected to cancel out in Eq. (38). Indeed, the behavior of $\Lambda(t)$ observed in successful retrieval cases supports this scenario (figure 8 (a)).
Conversely, when retrieval fails, the correlation function $C(t,s)$ decays as |ts||t-s| increases, exhibiting stationary behavior C(t,s)=C(|ts|)C(t,s)=C(|t-s|) (figures 6 (c,d)). The feedback coefficients $\Lambda(t,s)$ remain small except for the first few steps of $t$, while their integrated form continues to evolve slowly (figure 8 (b)), eventually converging to a finite value at longer time scales (inset of figure 8 (b)). Figure 9 shows the time correlation of $f(a_{i}^{t})$, $Q(t,s)=N_{\rm DMFT}^{-1}\sum_{i=1}^{N}f(a_{i}^{t})f(a_{i}^{s})$. In contrast to the case of successful retrieval (figure 9 (a)), the time correlation $Q(t,s)$ decays rapidly as |ts||t-s| grows in the failed retrieval case (figures 9 (b)). These observations suggest that the system stays in a chaotic state.
@@
Figure 9: Time correlation $Q(t,s)=N_{\rm DMFT}^{-1}\sum_{i=1}^{N}f(a_{i}^{t})f(a_{i}^{s})$. (a): Successful retrieval. (b): Failed retrieval.
@@ IV.3 Connection to earlier studies
for each $x$. The derivation of Eqs. (42)–(46) is provided in Appendix C. The convergent value of the readout overlap is assessed as $M=N_{\rm DMFT}^{-1}\sum_{i=1}^{N_{\rm DMFT}}{\rm sign}(m+\phi_{i})$. Comparison of Eqs. (41) and (45) indicates that $U=\sigma_{\phi}^{-2}K^{*}$ holds. Eqs. (42)–(46) are equivalent to the self-consistent signal-to-noise analysis (SCSNA) developed for analog neural networks in Ref. [18], provided that the sample averages $N_{\rm DMFT}^{-1}\sum_{i=1}^{N_{\rm DMFT}}(\cdots)$ are replaced by Gaussian averages $\int d\phi\,\mathcal{N}(\phi|0,\sigma_{\phi}^{2})(\cdots)$, which is justified by the law of large numbers in the limit of $N_{\rm DMFT}\to\infty$. Similar results can also be found in [19].
Ref. [18] shows that SCSNA can accurately describe macroscopic quantities such as storage capacity and the local-field distribution as long as the function $g(x)$ is uniquely determined by Eq. (46). However, it is also reported that when multiple solutions of Eq. (46) emerge as the profile of $f(x)$ is gradually modified, the SCSNA predictions become unreliable. Indeed, in the present model, the convergent DMFT solution reveals that Eq. (46) admits multiple branches while $g(x)$ is a single valued function for a monotonic transfer function (figures 10 (a,b)) For each input $m+\phi_{i}$ in Eqs. (42)–(44), the appropriate branch of $g(m+\phi_{i})$ is determined by the entire history of the DMFT dynamics. This implies that DMFT is indispensable for examining the fixed-point structure of the retrieval dynamics when non-monotonic transfer functions are used.
Figure 10: Markers show input-output relations after 1000 iterations of Eq. (29), which practically correspond to the solutions $g(x)$ of Eq. (46). (a) Monotonic transfer function $\tanh(10x)$ with $\alpha=0.1$. (b) Non-monotonic transfer function given by Eq. (8) with $\alpha=0.3$. In both cases, the initial state was set to the target pattern $(1,1,\ldots,1)^{\top}$. In contrast to the monotonic case (a), multiple branches of $g(x)$ emerge in the non-monotonic case (b). Red dotted lines stand for the transfer functions.
@@ Acknowledgements.
Let us denote $(T+1)\times(T+1)$ matrices composed of $Q(t,s)$, $R(t,s)$, and $S(t,s)$ in Eqs. (21)–(23) by $Q$, $R$, and $S$, respectively. Using these quantities, we define a $2(T+1)\times 2(T+1)$ matrix as
@@
and similarly for $R(t,s)$ and $S(t,s)$ for $\forall{t},\forall{s}\in\{0,\ldots,T\}$, into the right-hand side of Eq. (18). The normalization property $\lim_{\tilde{\bm{\theta}}\to\bm{0}}Z(\tilde{\bm{\theta}})=1$ of Eq. (12) for the addition of any external inputs in Eq. (11) ensures $R=0$. Evaluating integrals over $Q$, $R$, and $S$ using the saddle point method yields
@@
$\times\!\exp\!\left(-\!\sum_{i=1}^{N}\!\sum_{t,s}\!\left[\tfrac{1}{2}\hat{R}(t,s)(ih_{i}^{t})(ih_{i}^{s})+\hat{S}(t,s)f(a_{i}^{t})(ih_{i}^{s})\right]\!\right)\!\left.\times\exp\!\left(i\sum_{t=0}^{T+1}\bm{\theta}^{t}\!\cdot\!\bm{a}^{t}\right)\right\}$ (58)
$=$ $\,\mathop{\rm tr}_{\bm{\sigma}^{0},\tilde{\bm{a}},\tilde{\bm{h}},\tilde{\bm{\phi}}}\!\left\{P(\bm{\sigma}^{0})\exp\!\left[-\!\sum_{t=0}^{T}\sum_{i=1}^{N}ih_{i}^{t}{\mathcal{L}}_{i}^{t}\right]\times\!\exp\!\left(i\sum_{t=0}^{T+1}\bm{\theta}^{t}\!\cdot\!\bm{a}^{t}\right)\right$. (59)
@@
$\mathbb{E}_{\tilde{\bm{\phi}}}\left[\frac{\partial f(a_{i}^{t})}{\partial\phi_{i}^{s}}\right]$ $=\int\prod_{u=0}^{T}d\phi_{i}^{u}\frac{\exp\!\left(-\tfrac{\gamma^{2}}{2}\!\sum_{t,s}\!\hat{R}^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)}{(2\pi)^{(T+1)/2}\!\det(\gamma^{-2}\hat{R})^{1/2}}\frac{\partial f(a_{i}^{t})}{\partial\phi_{i}^{s}}$ (67)
$=\int\prod_{u=0}^{T}d\phi_{i}^{u}\frac{\exp\!\left(-\tfrac{\gamma^{2}}{2}\!\sum_{t,s}\!\hat{R}^{-1}(t,s)\phi_{i}^{t}\phi_{i}^{s}\right)}{(2\pi)^{(T+1)/2}\!\det(\gamma^{-2}\hat{R})^{1/2}}f(a_{i}^{t})\left(\sum_{w}\phi_{i}^{w}\gamma^{2}\hat{R}^{-1}(w,s)\right)$ (68)
$=\mathbb{E}_{\tilde{\bm{\phi}}}\left[f(a_{i}^{t})\left(\sum_{w}\phi_{i}^{w}\gamma^{2}\hat{R}^{-1}(w,s)\right)\right],$ (69)